what is 40 feet to 140 feet as 8 feet is to 28 feet

Answer by oberobic(2304)

40 is to 140 can be viewed as
40/140 = 4/14
4/14 = 2/7
.
8/28 = 4/14
.
So, yes, 40 is the same proportion of 140 as 8 is of 28.
.
Done.